OpenStudy (anonymous):

PLEASE HELP!! Solve 2(x + 1) = 2x + 2. All real numbers No solution 0 1

OpenStudy (anonymous):

Can you distribute? Or do you know how to?

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

Just a quick question, what grade are you in and what math class?

OpenStudy (kendricklamar2014):

All real numbers are solutions

OpenStudy (anonymous):

Okay Kendrick, how is it help if all you do is give an answer?

OpenStudy (kendricklamar2014):

Give me a sec

OpenStudy (kendricklamar2014):

1. Simplify both sides 2x +2 = 2x + 2

OpenStudy (kendricklamar2014):

2. Subtract 2 from both sides. 2x + 2 -2x = 2x +2 -2x

OpenStudy (kendricklamar2014):

Thats equals 2 =2 Step 3: Subtract 2 from both sides 2-2 = 2-2

OpenStudy (kendricklamar2014):

the equals 0=0, so all real numbers are a solution

OpenStudy (anonymous):

Great job. This simply means that whatever you put in for x both equations will always equal each other.

OpenStudy (kendricklamar2014):

Exactly ^

OpenStudy (anonymous):

so the answer is all real numbers?

OpenStudy (kendricklamar2014):

Yes